OpenStudy (anonymous):

A class aquarium is in the same shape of a rectangular prism measuring 2ft x 1ft x 1.5ft. A large fish requires 212in^3 of space, and a small fish requires 114in^3. If the tank is to have exactly 30 fish, what is the maximum number of large fish the class can purchase?

OpenStudy (allank):

This problem can be solved by optimization. Are you familiar with that?

OpenStudy (allank):

Optimization is where you find the maximum/minimum of a function, by using its first derivative and stuff like that. Does that sound familiar? I can walk you through it.

OpenStudy (allank):

It's quite simple really.

OpenStudy (allank):

Wait...actually we can solve this by using simultaneous equations. Sorry about that.

OpenStudy (anonymous):

Can you walk me through it please? i don't think i know how to do "optomization" or "stimultaneous equations" :O

OpenStudy (allank):

Alright. We are going to use simultaneous equations. Now, letting the number of large fish be L and small fish be S, we can say that L+S=30, because that's the maximum number of fish that can be bought. right?

OpenStudy (anonymous):

yupp

OpenStudy (allank):

Nice. We can also say that 212L+114S=(2*1*1.5)=3, because that's the maximum volume of the aquarium. Right?

OpenStudy (anonymous):

yes

OpenStudy (allank):

So we have two simultaneous equations: L+S=30 212L+114S=3 Are you with me till here?

OpenStudy (anonymous):

yeah, so do i just get either L or S alone on the 1st equation and then plug it in on the 2nd equation?

OpenStudy (allank):

Yep.

OpenStudy (anonymous):

i got a decimal... L=8.25991

OpenStudy (allank):

If you did the math right, that should be okay, and you'll thus round down the answer to 8 fish.

OpenStudy (anonymous):

So do i plug L back into the equation or is 8 just the answer?

OpenStudy (allank):

8 is the answer.

OpenStudy (anonymous):

Thank you sooo much :D

OpenStudy (allank):

You're welcome :)